OpenStudy (anonymous):

Multiply and simplify?????? just double checking (9x^7/15x^5)(5x^2/45x^3)= x^4/15x^3???

OpenStudy (science0229):

\[\frac{ 9x ^{7} }{ 15x ^{5} }*\frac{ 5x ^{2} }{ 45x ^{3} }=\frac{ 3x ^{2} }{ 5 }*\frac{ 1 }{ 9x }=\frac{ 3x ^{2} }{ 45x }=\frac{ x }{ 9 }\]. I think you made a mistake somewhere in your calculation.